THE PURE ROTATIONAL SPECTRUM OF TIS (X^{3}Δ_{r}) IN ALL THREE SPIN COMPONENTS

\underline{R. L. PULLIAM}, and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of TiS in the X^{3}Δ_{r} ground state has been measured using millimeter-wave direct absorption techniques. This species was created by the reaction of hydrogen sulfide gas with titanium vapor produced in a Broida oven which was modified to withstand the extreme temperatures required to melt this metal. Eight transitions have been recorded, each consisting of three spin components. No lambda-doubling was observed, as expected for a Δ state. The data were fit with a Hunds case(a) Hamiltonian and spectroscopic constants have been determined. The results from this study, as well as a comparison to past optical data will be presented.